# Reconstruction of frequency-localized functions from pointwise samples via least squares and deep learning

### v1

@@ 3 Main results
$\gamma(\mathsf{w}):=\lceil\log_{2}(30\mathsf{w})\rceil,$ (3.1)
to be used in the ensuing statements. We begin in Subsection 3.1 with our least squares approximation result. Then, our PET is presented in Subsection 3.1. We conclude in Subsection 3.3 with two additional least squares corollaries.
@@ 3.1 Least squares approximation
Our first result provides an error bound for the least squares problem (2.8) of approximating a continuous function using Slepian basis functions indexed over a hyperbolic cross. A proof can be found in Subsection 5.2.
@@
Let $\beta,\delta\in(0,1)$, $d=1,2,3$, and $\Lambda=\Lambda^{\rm HC}_{n-1}$ be the $d$-dimensional hyperbolic cross of order $n$, where $n\in\mathbb{N}$ if $d=1,2$, and $n\geq 26$ if $d=3$. For $\mathsf{w}\geq 1$, let $\mathcal{S}_{\Lambda}$ be the linear span of $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ and let $\vec{y}_{1},\dots,\vec{y}_{m}$ be random sample points drawn independently from the uniform measure $\bf u$ on $[-1,1]^{d}$. If $m\geq c_{\delta}\kappa^{\star}\log(\texttt{\#}\Lambda/\beta)$, where
$\kappa^{\star}:=(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}\quad\text{ and }\quad c_{\delta}:=((1-\delta)\log(1-\delta)+\delta)^{-1},$ (3.2)
@@
Theorem 3.1 concludes that frequency-localized functions that are well-approximated by the $\mathsf{w}$-bandlimited Slepian basis can be accurately recovered from pointwise samples. The number of samples required scales as $\kappa^{\star}=(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}$, up to constants and a logarithmic factor. The corresponding $L^{2}_{\bf u}$-error is dominated by the $L^{\infty}_{\bf u}$ distance of $f$ from the given Slepian basis and the $\ell^{2}$-norm of the noise vector, up to multiplicative constants. In particular, $\mathsf{w}$-bandlimited functions that are exactly linear combinations of elements in $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ can be recovered with accuracy proportional to the noise level $\|\vec{e}\|_{2}$. The error estimate in Theorem 3.1 provides an example of a uniform recovery guarantee, meaning that a single draw of sample points is sufficient to recover any function $f\in\mathcal{C}([-1,1]^{d})$ with high probability up to the bound given in (3.3). Further recovery guarantees for least squares are presented in Subsection 3.3.
@@ 3.2 Practical existence theorem
Our second result presents a PET for the least squares approximation of continuous functions, with a complexity rate derived from Theorem 3.1. A proof is given in Subsection 5.4.
@@
Let $\varepsilon,\delta,\beta\in(0,1)$, $d=1,2,3$, and $\Lambda=\Lambda^{\rm HC}_{n-1}$ be the $d$-dimensional hyperbolic cross of order $n$, where $n\in\mathbb{N}$ if $d=1,2$, and $n\geq 26$ if $d=3$. Let $\mathsf{w}\geq 1$ and $B(d,n)$, $M(d,n)$ be as in (3.4), (3.5), respectively. Let $\vec{y}_{1},\dots,\vec{y}_{m}$ be independent sample points drawn from the uniform measure $\bf u$ on $[-1,1]^{d}$, with $m\geq c_{\delta}\kappa^{\star}\log(\texttt{\#}\Lambda/\beta)$, where $\kappa^{\star}$, $c_{\delta}$ are given in (3.2). Then for every $\varepsilon>0$ sufficiently small satisfying
@@
Similar to Theorem 3.1, Theorem 3.2 states that functions well-approximated by the Slepian system $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ can be accurately recovered via deep learning using a training set of pointwise samples with size proportional to $\kappa^{\star}=(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}$ (up a constant and a log factor). Theorem 3.2 assumes training to be performed by solving a least squares-type optimization problem (3.7) that closely resembles (2.8), operating over a class of NNs $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ with specific architectures, determined by $\Lambda$, $\mathsf{w}$, $\varepsilon$. The class $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ is defined in (5.73) and is composed by NNs whose first-to-second-last layers are explicitly constructed to emulate the Slepian basis and whose last layer is made of trainable weights, optimized through (3.7). Theorem 3.2 provides explicit bounds on the depth $L(\Psi^{\natural})$ and width $W(\Psi^{\natural})$ of NNs in the class $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$, while the dependence of $N_{\star}$ on $\Lambda$, $\mathsf{w}$, $\varepsilon$ is detailed later in the proof of Theorem 3.2, particularly in (5.38). The error bound (3.2) of Theorem 3.2 is similar to (3.3) in Theorem 3.1, so that besides minor changes in the constant factors, the main difference between (3.2) and (3.3) is the presence of an extra term in the best approximation error. The impact of this extra term is mild, though, as it scales linearly with an auxiliary parameter $\varepsilon\in(0,1)$ that can be made arbitrarily close to 00 at the price of increasing the architecture bounds proportionally to $\log(1/\varepsilon)$.
@@ 3.3 Discussion on further recovery guarantees for least squares
In the same setting of Theorem 3.1, let $\beta,\delta\in(0,1)$ and $f\in\mathcal{C}([-1,1]^{d})$. Assume that the number of samples $m$ satisfies the condition $m\geq c\kappa^{\star}\log(2\texttt{\#}\Lambda/\beta),$ where $c>0$ is a universal constant, and $\kappa^{\star}$ is defined as in Theorem 3.1. Then for every noise vector $e\in\mathbb{C}^{m}$, with probability at least $1-\beta$, the least squares solution $f^{\natural}$ obtained in (2.8) satisfies
@@
This result can be obtained from Theorem 5.7 and Proposition 5.8 by applying the same argument as in [3, Corollary 5.10]. The error bound now contains the best approximation error of $f$ with respect to the $L^{2}_{\bf u}$-norm as desired. However, as opposed to Theorem 3.1, Corollary 3.3 is a nonuniform recovery guarantee, i.e., the error bound holds with probability $1-\beta$ for a fixed $f$. In other words, one single draw of sample points is not sufficient for the validity of the recovery guarantee for all $f\in\mathcal{C}([-1,1]^{d})$ simultaneously. Another limitation of Corollary 3.3 is the poor scaling of the error bound with respect to the failure probability $\beta$. The following result addresses these limitations by providing a recovery guarantee in expectation.
@@
This result can obtained from Theorem 5.7 and Proposition 5.8 by applying the same argument as in [3, Corollary 5.11] (originally proposed in [15, Theorem 2]). Notably, the poor scaling with respect to $\beta$ in the error bound of Corollary 3.3 is not there anymore. However, this comes at the price of having a recovery guarantee in expectation, as opposed to high probability, while also requiring a priori knowledge of a constant $L$ such that $\|f\|_{L_{{\bf u}}^{2}([-1,1]^{d})}\leq L$.
We conclude with a brief remark that the proof of Theorem 3.2 can also be modified to incorporate $L^{2}_{\bf u}$-$L^{2}_{\bf u}$ recovery guarantees by employing Corollary 3.3 or 3.4 in place of Theorem 3.1. For the sake of conciseness, we omit the statements of the resulting corollaries.
@@ 4 Numerical experiments
Benchmark functions.

### v2

@@ 3 Main results
$\gamma(\mathsf{w}):=\lceil\log_{2}(24\mathsf{w})\rceil,$ (3.1)
to be used in the ensuing statements. We begin in Section 3.1 with our least squares approximation result. Then, our PET is presented in Section 3.1. We conclude in Section 3.3 with two additional least squares corollaries.
@@ 3.1 Least squares approximation
Our first result provides an error bound for the least squares problem (2.8) of approximating a continuous function using Slepian basis functions indexed over a hyperbolic cross. A proof can be found in Section 5.3.
@@
Let $\beta,\delta\in(0,1)$, $d=1,2,3$, and $\Lambda=\Lambda^{\rm HC}_{n-1}$ be the $d$-dimensional hyperbolic cross of order $n$, where $n\in\mathbb{N}$ if $d=1,2$, and $n\geq 26$ if $d=3$. For $\mathsf{w}\geq 1$, let $\mathcal{S}_{\Lambda}$ be the linear span of $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ and let $\vec{y}_{1},\dots,\vec{y}_{m}$ be random sample points drawn independently from the uniform measure $\bf u$ on $1,1]^{d}$. If
$m\geq((1-\delta)\log(1-\delta)+\delta)^{-1}(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}\log(\texttt{\#}\Lambda/\beta),$ (3.2)
@@
The error estimate in Theorem 3.1 provides an example of a uniform recovery guarantee, meaning that a single draw of sample points is sufficient to recover any function $f\in\mathcal{C}([-1,1]^{d})$ with high probability up to the bound given in (3.3). The number of samples required in (3.2) scales as $(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}$, up to a constant and a logarithmic factor. The corresponding $L^{2}_{\bf u}$-error is dominated by the $L^{\infty}$-distance of $f$ from the given Slepian basis and the $\ell^{2}$-norm of the noise vector, up to multiplicative constants. Consequently, while Theorem 3.1 applies for any continuous function $f$, it is most informative for frequency-localized functions that are well-approximated by the Slepian span of $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$. For these functions, the approximation error term $\inf_{g\in\mathcal{S}_{\Lambda}}\|f-g\|_{L^{\infty}([-1,1]^{d})}$ is minimal, making the bound in (3.3) conclusive, and ensuring that pointwise samples suffice for accurate recovery. Particularly, $\mathsf{w}$-bandlimited functions that are exactly linear combinations of elements in $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ can be recovered with accuracy proportional to the noise level $\|\vec{e}\|_{2}$.
Further recovery guarantees for least squares are presented in Section 3.3.
Remark 3.2 (Comparison with polynomial least squares approximation theory).
Most of the literature on function approximation via least squares from random samples employs orthogonal polynomials (e.g., Legendre or Chebyshev) as the approximation basis—see the key contributions [chkifa2015discrete, cohen2013stability, cohen2017optimal, migliorati2014analysis, migliorati2013polynomial] and the surveys [adcock2022sparse, cohen2015approximation] for further references. The proof of Theorem 3.1 follows the same strategy used in the polynomial least squares literature, which relies on the study of the so-called Christoffel function—a key topic of Section 5.3. These polynomial least squares approximation results yield sample complexity bounds that are, at first glance, more favorable than (3.2). For example, under the same assumptions as in Theorem 3.1 it can be shown that (up to a logarithmic factor) a number of random samples $m$ scaling like $(\texttt{\#}\Lambda)^{2}$ is sufficient to recover the best approximation $L^{\infty}$-error of $f$ with respect to Legendre polynomials. A similar result holds for Chebyshev polynomials and random samples distributed according to the arcsine measure with a scaling proportional to $(\texttt{\#}\Lambda)^{\log_{3}(2)}\approx(\texttt{\#}\Lambda)^{1.58}$ (up to a logarithmic factor). Moreover, resorting to weighted least squares and picking the random samples according to the so-called Christoffel sampling, one obtains a log-linear sampling complexity (scaling linearly in $\texttt{\#}\Lambda$ up to a logarithmic factor) [cohen2017optimal].
Despite their seemingly more favorable sampling complexity, recovery guarantees for orthogonal polynomials are not directly comparable to Theorem 3.1. Indeed, the right-hand side of the error bound (3.3) (which has an analogous form in the orthogonal polynomial setting—see, e.g., [adcock2022sparse, Corollary 5.9]) contains the best $L^{\infty}$-approximation error of $f$, which depends on the choice of basis. Consequently, the relevant question becomes: which basis is more suitable to efficiently approximate $f$ (e.g., with a potentially smaller set $\Lambda$)? To the best of our knowledge, this approximation-theoretical question has no definitive answer in the literature. An interesting theoretical comparison is carried out in, e.g., [boyd2004prolate, Section 7], where the benefits of using the Slepian basis to approximate the simple analytic function $f(x)=\exp(-ikx)$ are shown asymptotically for $k\to\infty$. As we see in Section 4, there are functions (such as those considered in [boyd2004prolate]) for which least squares approximation from random samples is more efficient when employing Slepian basis than orthogonal polynomial bases.
@@ 3.2 Practical existence theorem
Our second result presents a PET for the least squares approximation of continuous functions, with a complexity rate derived from Theorem 3.1. A proof is given in Section 5.5.
@@
Let $\varepsilon,\delta,\beta\in(0,1)$, $d=1,2,3$, and $\Lambda=\Lambda^{\rm HC}_{n-1}$ be the $d$-dimensional hyperbolic cross of order $n$, where $n\in\mathbb{N}$ if $d=1,2$, and $n\geq 26$ if $d=3$. Let $\mathsf{w}\geq 1$ and $B(d,n)$, $M(d,n)$ be as in (3.4), (3.5), respectively. Let $\vec{y}_{1},\dots,\vec{y}_{m}$ be independent sample points drawn from the uniform measure $\bf u$ on $1,1]^{d}$, with
$m\geq((1-\delta)\log(1-\delta)+\delta)^{-1}(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}\log(\texttt{\#}\Lambda/\beta).$
Then for every $\varepsilon>0$ sufficiently small satisfying
@@
A few remarks are in order. First, similar to Theorem 3.1, Theorem 3.3 states that functions well-approximated by the Slepian span of $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ can be accurately recovered via deep learning using a training set of pointwise samples with size proportional to $(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}$, up a constant and a logarithmic factor. The theorem assumes training to be performed by solving a least squares-type optimization problem (3.7) that closely resembles (2.8), operating over a class of NNs $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ with specific architectures, determined by $\Lambda$, $\mathsf{w}$, and $\varepsilon$. The class $\mathcal{N}_{\Lambda,\mathsf{w},\varepsilon}$ is defined in (5.70) and is composed by NNs whose first-to-second-last layers are explicitly constructed to emulate the Slepian basis and whose last layer is made of trainable weights, optimized through (3.7).
Second, we briefly discuss the upper bounds on $L(\Psi^{\sharp})$ and $W(\Psi^{\sharp})$, i.e., the depth and width of the optimal network $\Psi^{\sharp}$, respectively. Note that since training occurs only in the last layer, these bounds are not directly tied to the computational difficulty of training $\Psi^{\sharp}$. However, they explicitly quantify how the network depth and width depend on $\Lambda$, $\mathsf{w}$, $\varepsilon$. In order to better understand these bounds, we unpack the definition of $N_{\star}=N_{\star}(\Lambda,\mathsf{w},\varepsilon)$. For a one-dimensional index set $\Lambda$, this is given in (5.35). For the special case $\Lambda=\Lambda^{\rm HC}_{n-1}$ where $d=1,2,3$, the dependence on $\Lambda,\mathsf{w},\varepsilon$ can be simplified to that on $\mathsf{w},\varepsilon$, since—as explained in the proof of Proposition 5.10—we have
$N_{\star}=\Big\lceil\max\Big\{2\lfloor e\mathsf{w}\rfloor+1,\frac{\log(3/(\varepsilon\mu_{\mathsf{w},n-1}))}{\log(3/2)}\Big\}\Big\rceil.$ (3.9)
Here, recall that $\mu_{\mathsf{w},n-1}$ denotes the $n$th-eigenvalue (2.4) of the operator $P_{\mathsf{w}}Q_{T}$. Moreover, for $\mathsf{w}\geq 1$, the following bounds, adapted from [karnik2021improved, Corollary 3], hold for a universal constant $C>0$:
[…]
From (3.9), (3.10), it follows that when $\Lambda=\Lambda^{\rm HC}_{n-1}$ with $d=1,2,3$, the quantity $N_{\star}$ grows linearly with respect to $\mathsf{w}$ and logarithmically with $1/\varepsilon$. Applying this to Theorem 3.3 yields a network depth and width that scale at most polynomially with $\mathsf{w}$ and logarithmically with $1/\varepsilon$.
Third, the error bound (3.3) in Theorem 3.3 is similar to (3.3) in Theorem 3.1, so that besides minor changes in the constant factors, the main difference between (3.3) and (3.3) is the presence of an extra term in the best approximation error. The impact of this extra term is mild, though, as it scales linearly with an auxiliary parameter $\varepsilon\in(0,1)$ that can be made arbitrarily close to 0 at the price of increasing the architecture bounds proportionally to $\log(1/\varepsilon)$.
Finally, we note that Theorem 3.3 demonstrates the existence of a class of neural networks that, when trained, can attain performance nearly matching least squares approximation in the Slepian basis. Owing to the nature of its proof (see Section 5.5), however, the theorem does not assert that such networks can outperform the Slepian basis—though this may occur empirically. Moreover, we stress that the main purpose of Theorem 3.3 is of theoretical nature (i.e., it provides an improved version of a universal approximation theorem), and it is not intended to propose a new algorithmic recipe.
@@ 3.3 Discussion on further recovery guarantees for least squares
In the same setting of Theorem 3.1, let $\beta,\delta\in(0,1)$ and $f\in\mathcal{C}([-1,1]^{d})$. Assume that the number of samples $m$ satisfies the condition $m\geq c(2\texttt{\#}\Lambda)^{2\gamma(\mathsf{w})}\log(2\texttt{\#}\Lambda/\beta)$, where $c>0$ is a universal constant. Then for every noise vector $e\in\mathbb{C}^{m}$, with probability at least $1-\beta$, the least squares solution $f^{\natural}$ obtained in (2.8) satisfies
@@
This result can be obtained from Theorem 5.7 and Proposition 5.8 by applying the same argument as in [adcock2022sparse, Corollary 5.10]. The error bound now contains the best approximation error of $f$ with respect to the $L^{2}_{\bf u}$-norm as desired. However, as opposed to Theorem 3.1, Corollary 3.4 is a nonuniform recovery guarantee, i.e., the error bound holds with probability $1-\beta$ for a fixed $f$. In other words, one single draw of sample points is not sufficient for the validity of the recovery guarantee for all $f\in\mathcal{C}([-1,1]^{d})$ simultaneously. Another limitation of Corollary 3.4 is the poor scaling of the error bound with respect to the failure probability $\beta$. The following result addresses these limitations by providing a recovery guarantee in expectation.
@@
This result can be obtained from Theorem 5.7 and Proposition 5.8 by applying the same argument as in [adcock2022sparse, Corollary 5.11] (originally proposed in [cohen2013stability, Theorem 2]). Notably, the poor scaling with respect to $\beta$ in the error bound of Corollary 3.4 is not there anymore. However, this comes at the price of having a recovery guarantee in expectation, as opposed to high probability, while also requiring a priori knowledge of a constant $L$ such that $\|f\|_{L_{{\bf u}}^{2}([-1,1]^{d})}\leq L$.
We conclude with a brief remark that the proof of Theorem 3.3 can also be modified to incorporate $L^{2}_{\bf u}$-$L^{2}_{\bf u}$ recovery guarantees by employing Corollary 3.4 or 3.5 in place of Theorem 3.1. For the sake of conciseness, we omit the statements of the resulting corollaries.
@@ 4 Numerical experiments
Training set, test set, and error metric.
For a given target function $f$ that we seek to approximate using either Least Squares or Deep Learning, we generate data sets of the form $\{(x_{j},f(x_{j})\}_{j=1}^{m}$ with $x_{j}$ drawn independently from the uniform distribution on $1,1]^{d}$. For a set of $m$ training points, we also consider a test set $\{x_{j}^{\text{test}}\}_{j=1}^{m_{\text{test}}}$ of size $m_{\text{test}}=0.2m$, sampled independently according to the uniform distribution on $1,1]^{d}$. The corresponding test error is given by the Root Mean Square Error (RMSE), defined as
where $f$ is the target function, and $\tilde{f}$ is the approximation obtained from either method.
Slepian basis functions.
To construct one-dimensional Slepian basis functions $\varphi_{\mathsf{w},j}$, we employ the Python package scipy.special [2020SciPy-NMeth], which is based on the angular solutions of the Helmholtz wave equation of the first kind. In fact, each function $\varphi_{\mathsf{w},j}$ is a scalar multiple of a corresponding angular solution [moore2004prolate]. To construct two-dimensional Slepian basis functions, we utilize the tensor-product definition (2.7) $\varphi_{\mathsf{w},\vec{\nu}}=\varphi_{\mathsf{w},\nu_{1}}\otimes\varphi_{\mathsf{w},\nu_{2}}$. Finally, in both cases, we take the index set to be the hyperbolic cross $\Lambda^{\text{HC}}_{n-1}$ (Definition 2.3) for various orders $n\in\mathbb{N}$.
4.1 The Slepian basis vs. orthogonal polynomials
To begin, we illustrate how the Least Squares method with the Slepian basis can outperform other recurrent orthogonal bases. More specifically, we compare the Slepian basis with the Chebyshev and Legendre orthogonal polynomial bases. Recall that Chebyshev polynomials (of the first kind) are defined on the interval [1,1][-1,1] as $T_{k}(x)=\cos(k\arccos x)$, for $k\in\mathbb{N}$. For a definition of Legendre polynomials, see Equation (5.33) in Section 5.4, where they are used to prove Theorem 3.3. For more information about these families of orthogonal polynomials, we refer to, e.g., [adcock2022sparse, Section 2.2]. To perform our analysis, we consider the following functions. In one dimension, we take the Gaussian
$g_{1}(x):=e^{-\pi x^{2}},\qquad\forall x\in[-1,1],$
and the complex exponential
$g_{2}(x):=e^{4\pi ix},\qquad\forall x\in[-1,1].$
In two dimensions, we consider
$g_{3}(x,y):=e^{-\pi(x^{2}+y^{2}+0.2xy)},\qquad\forall(x,y)\in[-1,1]^{2}.$
Note that $g_{3}$ is non-separable, meaning it is not simply the product of two one-dimensional functions, thus making it purely two-dimensional. For each of these functions, we run two types of experiments. First, we analyze how the approximation error behaves when the number of samples varies while keeping the basis fixed as $\Lambda^{\text{HC}}_{n-1}$ (note that in the one-dimensional case $\Lambda^{\text{HC}}_{n-1}=\{0,\ldots,n-1\}$ corresponds to $n$ basis elements). For each of the functions $g_{1},g_{2},g_{3}$, the sample size $m$ takes values from $10$ up to $10000$. For every experiment using a set of $m$ training points, we also consider a test set $\{x_{j}^{\text{test}}\}_{j=1}^{m_{\text{test}}}$ of size $m_{\text{test}}=1000$, sampled independently from the uniform distribution on $1,1]^{d}$. Second, we observe the behavior of the error when the number of samples is fixed at $m=1000$ and the order $n$ varies.
Figure 2: Comparison of Least Squares approximations using different bases for the function $g_{1}(x)=e^{-\pi x^{2}}$ from $m$ random samples. The left panel shows results using $n=10$ basis functions and varying the number of samples, while the right panel shows results when the number of samples is fixed at $m=1000$ and the order parameter $n$ changes. Note that the results for Chebyshev and Legendre polynomials can be hard to distinguish because they are almost identical.
Figure 3: Comparison of least squares approximations using different bases for the function $g_{2}(x)=e^{4i\pi x}$ from $m$ random samples. The left panel shows results using $n=20$ basis functions and varying the number of samples, while the right panel shows results when the number of samples is fixed at $m=1000$ and the order parameter $n$ changes. Note that the results for Chebyshev and Legendre polynomials can be hard to distinguish because they are almost identical.
One-dimensional results.
Figures 2 and 3 display the RMSEs for the functions $g_{1}$ and $g_{2}$, respectively. For moderate order $n$, using the Slepian basis for Least Squares reconstruction outperforms the Chebyshev and Legendre bases. (Note, when $d=1$ we simply have $\Lambda^{\text{HC}}_{n-1}=\{0,1,\ldots n-1\}$, hence $n$ is also the number of basis functions.) In particular, when the number of samples is large enough, the RMSE for the Slepian basis is between one and two orders of magnitude smaller than for the two polynomial bases. When the order $n=10$ is fixed, the RMSE converges as the number of samples reaches $m=1000$ and remains stable. At $m=1000$ samples, the RMSE decreases as the number of parameters increases. Reconstructions with the Chebyshev and Legendre bases yield very similar, though not identical, RMSEs, which are difficult to distinguish on the logarithmic scale.
Two-dimensional results.
Figure 4 shows that the Slepian basis is superior for reconstructing $g_{3}$.
Figure 4: Comparison of Least Squares approximations using different bases for the function $g_{3}(x)=e^{-\pi(x^{2}+y^{2}+0.2xy)}$ from random samples. The left panel shows results using a hyperbolic cross $\Lambda^{\text{HC}}_{n-1}$ with $n=10$ and varying the number of samples, while the right panel shows results when the number of samples is fixed at $m=1000$ and the order parameter $n$ changes. Note that the results for Chebyshev and Legendre polynomials can be hard to distinguish because they are almost identical.
In general, the RMSEs obtained in this case are several orders of magnitude higher than those in the one-dimensional case, reflecting the increased complexity of the approximation problem. In the left-hand plot, we consider $n=10$, which entails a large number of tensorized Slepian functions within the hyperbolic cross $\Lambda^{\text{HC}}_{n-1}$ (for a bound on its cardinality in all dimensions, see for example Lemma 6.1). As in the one-dimensional experiments, the RMSE converges as the number of samples increases, for both the Slepian and the polynomial bases. Likewise, the right-hand plot, where the number of samples is fixed at $m=1000$, shows that the RMSE decreases as the number of parameters increases. In general, the benefits of Slepian over orthogonal functions appear more evident for $g_{3}$ and $g_{1}$ than $g_{2}$.
4.2 Comparisons of different initializations for Deep Learning
Additionally, we conduct a set of experiments to compare different neural network initialization strategies. The main purpose of this section is to show that a novel Slepian-based initialization inspired by Theorem 3.3 can lead to more accurate approximations with respect to other standard random initializations strategies for Deep Learning. In particular, we train networks with architecture $(1,1000,1000,10,1)$, where, in general, an architecture $(n_{0},\ldots,n_{L+1})$ denotes a feedforward neural network with input layer of dimension $n_{0}$, output layer of dimension $n_{L+1}$ and $L$ hidden layers of size $n_{1},\ldots,n_{L}$, respectively. In our experiments, we consider the function
$f_{1}(x)=\cos(10x)e^{-\pi x^{2}},\qquad\forall x\in[-1,1].$ (4.1)
This one-dimensional function, defined as the product of a trigonometric function and a Gaussian, was considered in [moore2004prolate] as a simple example in which the Slepian basis achieves better approximation accuracy than the standard Fourier basis. In all cases, networks are trained using Adam [diederik2015adam] with an exponential decay rate of $0.85$.
We start by describing the proposed Slepian-based initialization strategy. Inspired by the decomposition of a frequency-localized function $f$ with respect to the Slepian basis,
$f(x)\approx\sum_{j=0}^{J}\alpha_{j}\varphi_{\mathsf{w},j},$ (4.2)
and by the construction of Theorem 3.3 (see Section 5.4), we first approximate each Slepian basis function $\varphi_{\mathsf{w},j}$ using distinct neural networks $\tilde{\varphi}_{\mathsf{w},j}$ with architecture $(1,100,100,1)$. Each network is trained on 10,00010{,}000 samples $(\vec{x}_{i},\varphi_{\mathsf{w},j}(\vec{x}_{i})),i=1,\dots,10000$, where $\vec{x}_{i}\in[-1,1]^{d}$ are drawn uniformly at random and for 1000 epochs. We train networks for the first ten Slepian basis functions ($j=0,1,\dots,9$). The test errors for these networks are reported in Table 1.
Index $j$ Training error Test error Index $j$ Training error Test error
0 $9.320\times 10^{-8}$ $8.837\times 10^{-8}$ 5 $1.571\times 10^{-4}$ $1.960\times 10^{-4}$
1 $4.365\times 10^{-7}$ $3.578\times 10^{-7}$ 6 $3.448\times 10^{-3}$ $3.270\times 10^{-3}$
2 $8.744\times 10^{-7}$ $7.630\times 10^{-7}$ 7 $2.218\times 10^{-4}$ $2.430\times 10^{-4}$
3 $2.680\times 10^{-7}$ $2.766\times 10^{-7}$ 8 $1.795\times 10^{-5}$ $1.845\times 10^{-5}$
4 $1.877\times 10^{-5}$ $1.921\times 10^{-5}$ 9 $307\times 10^{-4} 1.1$$138\times 10^{-4}$
Table 1: Test and training errors for networks $\tilde{\varphi}_{\mathsf{w},j}$ with architecture $(1,100,100,1)$ approximating the Slepian basis functions $\varphi_{\mathsf{w},j}$.
We observe that the approximation error is small for low-index basis functions and increases with $j$; in particular, the functions with indices $j=6,7$ yield the largest errors. Once each Slepian basis function is approximated, we combine the trained networks according to Equation (4.2) and determine the coefficients $\alpha_{j}$ via least-squares fitting using 10,00010{,}000 new random samples $(\vec{x}_{i},f(\vec{x}_{i}))$. The networks and weights are then concatenated into a single network $\tilde{\Phi}_{1}$ of architecture $(1,100\times 10,100\times 10,10,1)=(1,1000,1000,10,1)$ that realizes the corresponding linear combination. This construction is analogous to the neural network developed in Section 5.5 for the proof of Theorem 3.3. The resulting network $\Phi_{1}$ is subsequently used to initialize a network with identical architecture, denoted $\tilde{\Phi}_{1}$. Note that the most expensive part of this construction is training the networks $\tilde{\varphi}_{\mathsf{w},j}$. However, this computation is independent of the particular function $f$ that one wants to approximate. Hence, it can be considered as an offline pre-computing step. On the other hand, the least square fitting step is function dependent but can be solved quite efficiently.
In parallel, we train three additional networks with the same $(1,1000,1000,10,1)$ architecture but different initialization schemes, i.e.,
  • $\Phi_{2}$ with standard normal initialization (mean 0, standard deviation 0.1),
  • $\Phi_{3}$ with He initialization [he2015delving],
  • $\Phi_{4}$ with Glorot initialization [glorot2010understanding].
The corresponding training and test errors for $\Phi_{1}$, $\Phi_{2}$, $\Phi_{3}$, and $\Phi_{4}$ are shown in Figure 5.
Figure 5: Training and test errors for neural networks with architecture $(1,1000,1000,10,1)$ approximating the function $f_{1}(x)=\cos(10x)e^{-\pi x^{2}}$ using different initialization strategies. The left panel shows the training error and the right panel shows the test error.
Our proposed initialization ($\Phi_{1}$) outperforms all others, achieving errors on the order of $10^{-6}$ for both training and test sets (considering the best epoch among those shown). Glorot initialization yields errors around $10^{-2}$, while the standard normal and He initializations result in errors of approximately $10^{-1}$. Across all initialization methods, training and test errors converge within a few epochs.
4.3 Least Squares vs. Deep Learning